If 15.0 L of hydrogen chloride gas reacts with 48.0 L of oxygen gas, what is the volume of chlorine gas produced? Assume all gases are at the same temperature and pressure.

HCl(g)+O2(g)→Cl2(g)+H2O(g)
1
Expert's answer
2020-12-01T09:31:29-0500

If all gases are under same T and p, then gas volume is directly proportional to its amount.

4HCl + O2 = 2Cl2 + 2H2O

We have the excess of oxygen so we calculate everything from hydrogen chloride volume.

V(Cl2)=1/2V(HCl) = 15L/2 = 7.5L (stoichiometric coefficients)